how do you write this equation in slope intercept form ?
slope 3/4 passes through the origin ?

Answer by mananth(16946)

slope = 3/4
it is passing through origin
(0,0)

y= mx +b is the equation of the line with y intercept = b
since the line passes through origin b=0
y= mx
m=3/2
y= 3/4 x is the equation